OpenStudy (anonymous):

Find and simplify f(x)=6-7x^2 for: f(x+h)-f(x)

OpenStudy (anonymous):

\( f(x+h)=6-7(x+h)^2 \) and \( f(x)=6-7x^2 \) then \( f(x+h)-f(x)=6-7(x+h)^2-(6-7x^2)=6-7(x+h)^2-6+7x^2=7x^2-7(x+h)^2\) can u continue ?

OpenStudy (anonymous):

yeah i got to 6-7x^2+14xh+7h^2-6-7x^2 but not sure how to simplify after that

OpenStudy (anonymous):

-12+14xh-7h^2 would that be right?

OpenStudy (anonymous):

6-7x^2-14xh-7h^2-6+7x^2 -- -- -- compare with your answer and be carefull about signs

OpenStudy (anonymous):

it would be -14xh-7h^2

OpenStudy (anonymous):

or -14xh-7h^2?

OpenStudy (anonymous):

leme try it in my hw

OpenStudy (anonymous):

yes thats right try to work it out